Adult tickets for a play cost $25 each and the childrens tickets cost $15 each. The total receipts were $7,200 and the total attendenct was 400. How many adults and how many children attended?
Answer by checkley77(12844)

a=c=400 or a=400-c
25a+15c=7,200
25(400-c)+15c=7,200
10,000-25c+15c=7,200-10,000
-10c=-2,800
c=-2,800/-10
c=280 ans. for the number of children tickets sold.
400-280=120 adult tickets were sold.
Proof:
25*120+15*280=7,200
3,000+4,200=7,200
7,200=7,200